# Yoga for Fourier–Mukai partnership

### v1

@@ 1. Introduction
1.1. Setting the stage
The context of our work starts with Mukai [Muk81]. Loc. cit. introduced the notion of integral transforms to study triangulated equivalences between derived categories associated with abelian varieties. Over time, integral transforms have played a central role in the interplay between derived categories and algebraic geometry (e.g. [Bri02, Kaw02]), with further applications in string theory [AHR05, ASS05].
Much of this progress has focused on smooth varieties, where geometric results rely on categorical methods. However, extending these methods to singular varieties remains challenging, partly because bounded coherent complexes are typically not perfect. Thus, an important task is to develop new results or techniques that improve our understanding of integral transforms in such singular settings.
@@ 1.1. Setting the stage
In our work, we reveal additional depth to this perspective of exploiting special fibers. Specifically, we study the behavior of integral transforms under changes of the base scheme in more general settings that includes local algebra. Moreover, our approach naturally extends several results in the literature to arbitrary fields, and at the same time yields new insight for singularities of fibrations.
1.2. Preservation of adjointness
@@
For those concerned with equivalences, this allows the study of integral transforms to be reduced to their behavior on these subcategories. Whether an integral transform restricts to these smaller categories is now well-understood [Riz17, Bal09, DLM25]. These results shed light to the existence of left or right adjoints to integral transforms and the conditions under which such adjoints also restrict to these subcategories.
@@
1.3. Fully faithfulness & equivalences
Now, we have a clearer understanding of how integral transforms behave under base change. Particularly, regarding the preservation of adjointness and the restriction to smaller categories. Next, we shift gears to studying fully faithfulness and equivalences under base change; specifically, the ascent and descent of these properties. Before doing so, we review the relevant literature a bit more.
As mentioned earlier, the work of [HLS07, HLS09] studied these questions for fibrations whose base is of finite type over an algebraically closed field. A key ingredient in loc. cit. was to look at the behavior on special fibers. More recently, such questions have also been studied for smooth proper morphisms over a Noetherian base scheme (see [Kur24, Proposition/Definition 2.3]). However, smoothness constraints can restrict the examples which one might be interested in, e.g. fibrations admitting singular fibers.
A result of Orlov addresses similar questions in the case of smooth varieties over a field. Specifically, for smooth projective varieties over a field, an integral transform is fully faithful or an equivalence if, and only if, the induced integral transform satisfies the same property after any base field extension [Orl02, Lemma 2.12]. The strategy of loc. cit. leverages ‘canonical morphisms’ between the kernels of integral transforms that describe the (co)units of adjunction. However, it remains unclear whether such canonical morphisms exist in the singular setting. This uncertainty is another source of motivation for our work, as we aimed to bypass the lack of these tools in singular cases using different techniques.
The brings attention to the following, which summarizes Section˜4.
@@ 1.3. Fully faithfulness & equivalences
1.4. Affine variations
Naturally, we sought analogs of Theorem˜1.2 in which $t$ need not be locally of finite tor-dimension. Ideally, such a framework would include morphisms from special fibers as in [HLS07, HLS09]. Accordingly, we work in the setting of affine morphisms to the base, while imposing a mild restriction on one of the $Y_{i}\to S$.
The following result is a slight generalization of [HLS09, Proposition 2.15]. In particular, we perform base change along affine morphisms rather than closed immersions of closed points. Note that a key difference between what follows and Sections˜3 and 4 is that we impose one of the $Y_{i}\to S$ be locally projective (as opposed to proper like our earlier results).
Theorem 1.4 (see Corollary˜5.5).
Consider affine morphism $t\colon T\to S$ to a semiseparated Noetherian scheme. Suppose $Y_{1}\to S$ is flat, locally projective and $Y_{2}\to S$ is flat, proper. Denote by $t^{\prime}\colon Y_{1}\times_{S}Y_{2}\times_{S}T\to Y_{1}\times_{S}Y_{2}$ the natural morphism. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over each $Y_{i}$. If $\Phi_{K}$ restricts to a fully faithful functor (resp. an equivalence) on $D^{b}_{\operatorname{coh}}$, then so does $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$.
It is worthwhile to note that Theorem˜1.4 yields an independent proof of the ascending full faithfulness or equivalences in [HLS09, Proposition 2.15]. To do so, we show in suitable cases that the unit of an adjoint pair of integral transforms is preserved under base change when the left adjoint is fully faithful (see Proposition˜5.3 and Proposition˜4.2). However, it is not clear to us whether this remains true when the left adjoint is not fully faithful a priori. This condition seems necessary (at least to us) to establish the ‘descent’ part of [HLS09, Proposition 2.15] in the case of special fibers.
1.5. Applications
@@
To conclude our work, we discuss the ramifications of Theorems˜1.2 and 1.4. Our results extend the literature on schemes of finite type over algebraically closed fields to the more general setting of arbitrary fields. They reveal an interplay between the intuition of [HLS07, HLS09] concerning special fibers and our work on faithfully flat descent along affine morphisms. In particular, certain singularities are shown to be derived invariants for fibrations over Noetherian schemes, broadening the scope to include arithmetically flavored situations of potential interest for future research.
@@
For a pair of geometrically integral projective curves over a field $k$ of characteristic zero, Fourier–Mukai partnership implies they are $k$-forms (see Proposition˜6.1). Here, a $k$-form means there exists a finite extension $L/k$ such that their base changes along $L/k$ become isomorphic. The result holds for singular curves and enriches algebraically closed case of [Spe23]. In Example˜6.2, we exhibit a pair of non-isomorphic elliptic curves over $\mathbb{Q}$ that are nevertheless $\mathbb{Q}$-forms. For fibrations of relative dimension one over a Noetherian $\mathbb{Q}$-scheme, Fourier–Mukai partnership forces the genera of special fibers to coincide (see Corollary˜6.3).
@@
Smoothness for proper schemes over an arbitrary field is likewise invariant under Fourier–Mukai partnership (see Proposition˜6.4). This subtlety arises especially over imperfect fields. Since regularity is characterized by $D^{b}_{\operatorname{coh}}=\operatorname{Perf}$, one might expect this condition to persist after a change of field extensions, but there exist regular projective curves that become singular after such a base change (see e.g. [Kol11, Remark 16]).
@@
In the setting of projective fibrations, smoothness of the total space is also a derived invariant (see Corollary˜6.5). Finally, generalizing [HLS09], we show that the Cohen–Macaulay and Gorenstein properties of a morphism remain invariant under Fourier–Mukai partnership for Noetherian base schemes (see Proposition˜6.6).
@@
At times if $X$ is affine, we abuse notation and write $D(R):=D_{\operatorname{qc}}(X)$ where $R:=H^{0}(X,\mathcal{O}_{X})$ are the global sections; similar conventions will occur for the other categories.
@@ 2.1. Perfect & pseudocoherent complexes
We say $P\in D_{\operatorname{qc}}(X)$ is perfect if it is affine locally isomorphic to a bounded complex of locally free sheaves of finite rank. Denote by $\operatorname{Perf}(X)$ for the full subcategory of perfect complexes. Also, given an integer $N$, we say $E\in D_{\operatorname{qc}}(X)$ is $N$-pseudocoherent if locally there is a morphism $P\to E$ from some $P\in\operatorname{Perf}(X)$ such that the induced morphism on cohomology sheaves $\mathcal{H}^{n}(P)\to\mathcal{H}^{n}(E)$ is an isomorphism for $n>N$ and is surjective for $n=N$. We say $E\in D_{\operatorname{qc}}(X)$ is pseudocoherent if it is $N$-pseudocoherent for every $N\in\mathbb{Z}$. In general, since $X$ is a scheme, $E\in D_{\operatorname{qc}}(X)$ is pseudocoherent if and only if locally on $X$ there is a quasi-isomorphism $P\to E$, with $P$ a bounded above complex of finite-rank free modules [GW23, Corollary 22.47]. As $X$ is Noetherian, $E\in D_{\operatorname{qc}}(X)$ is pseudocoherent if, and only if, $\mathcal{H}^{n}(E)$ is coherent for all $n$ and vanishes for $n\gg 0$ (see e.g. [Sta25, Tag 08E8]). The subcategory of pseudocoherent complexes of $D_{\operatorname{qc}}(X)$ coincides with $D^{-}_{\operatorname{coh}}(X)$ (i.e. the full subcategory of complexes with bounded above and coherent cohomology). Consider a morphism $f\colon Y\to X$ of Noetherian schemes. Then one has $\mathbf{L}f^{\ast}D^{-}_{\operatorname{coh}}(X)\subseteq D^{-}_{\operatorname{coh}}(Y)$ (see e.g. [Sta25, Tag 08H4]). We will use this fact liberally.
@@ 2.3. Generation for triangulated categories
Let $\mathcal{T}$ be a triangulated category with shift functor $1]\colon\mathcal{T}\to\mathcal{T}$. Fix $\mathcal{S}\subseteq\mathcal{T}$. Denote by $\operatorname{add}(\mathcal{S})$ the strictly full (that is, closed under isomorphisms) subcategory of $\mathcal{T}$ consisting of all direct summands of objects of the form $\oplus_{n\in\mathbb{Z}}S_{n}^{\oplus r_{n}}[n]$ where $S_{n}\in\mathcal{S}$ for all $n$ and finitely many $r_{n}$ are nonzero. Inductively, let
@@
Now, we impose $\mathcal{T}$ be a compactly generated triangulated category. To clarify, we require that $\mathcal{T}$ admitting all small coproducts. Denote the collection of compact objects of $\mathcal{T}$ by $\mathcal{T}^{c}$. We say $G\in\mathcal{T}^{c}$ is a compact generator for $\mathcal{T}$ if for any $E\in\mathcal{T}$, one has $\operatorname{Hom}(G,E[n])=0$ for all $n\in\mathbb{Z}$ if, and only if, $E=0$. A well-known fact is that classical generators for $\mathcal{T}^{c}$ coincide with being compact generators for $\mathcal{T}$ (see e.g. [Sta25, Tag 09SR]).
@@
Let $X$ be a quasi-compact quasi-separated scheme. The collection of compact objects in $D_{\operatorname{qc}}(X)$ coincides with $\operatorname{Perf}(X)$. Moreover, $D_{\operatorname{qc}}(X)$ is compactly generated by a single object from $\operatorname{Perf}(X)$.
@@ 3. Preservation of adjointness
Consider ˜3.1 with $t$ a morphism locally of finite tor-dimension. Then $t_{1},t_{2},t^{\prime}$ are locally of finite tor-dimension. In particular, each face of Equation˜3.1 is a tor-independent square (in the sense of [Lip09, Definition 3.10.2]).
@@
This first claim is affine local on $S$. Indeed, if needed, we can find an affine open cover of $S$ such that the base changes of $t$ along the associated open immersion are of finite tor-dimension. Then, for each such open immersion of a component in the affine cover, we have a similar diagram to Equation˜3.1 (i.e. a ‘base change cube’). If we can show the desired properties on the morphisms arising in each diagram associated to the components of the affine open cover of $S$, then the desired claim follows. So, we may impose that $t$ be of finite tor-dimension.
By base change, we know that $g_{1},g_{2},f^{\prime}_{2},f^{\prime}_{1},g_{2}^{\prime},g_{1}^{\prime}$ are proper flat morphisms. Consider the fibered square in Equation˜3.1 induced from the base change of $f_{1}$ along $t$. It follows by [GW23, Lemma 23.27] that $t_{1}$ is of finite tor-dimension. A similar argument shows $t_{2}$ is of finite tor-dimension as well. In fact, if we work with the fibered square in Equation˜3.1 obtained by base change of $f_{2}^{\prime}$ along $t_{1}$, we see from similar reasoning that $t^{\prime}$ is locally of finite tor-dimension. This completes the proof for our first claim.
@@
Consider ˜3.1 with $t$ an affine morphism. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{1}$. Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor on $D^{b}_{\operatorname{coh}}$. Additionally, a similar statement holds for $\operatorname{Perf}$ if $K$ is relatively perfect over $Y_{2}$.
@@
We prove the case of $\Phi_{\mathbf{L}(t^{\prime})K}$ restricting to an exact functor on $\operatorname{Perf}$ because a similar argument applies for $D^{b}_{\operatorname{coh}}$. As $\Phi_{K}(\operatorname{Perf}(Y_{1}))\subseteq\operatorname{Perf}(Y_{2})$, [DLM25, Proposition 3.2] implies
@@
Consider ˜3.1. Let $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$. Then there is a natural isomorphism $\beta^{K}\colon\mathbf{L}t_{2}^{\ast}\circ\Phi_{K}\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\circ\mathbf{L}t_{1}^{\ast}$.
@@
This follows from the string of isomorphisms for each $E\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)$:
@@
However, we have an isomorphism in Equation˜3.1,
@@
$=t_{2}(\operatorname{supp}(\mathcal{H}^{j}(\mathbf{L}t_{2}^{\ast}\Phi_{K}(E)))).$
@@
Consider ˜3.1 where $t$ is an affine morphism locally of finite tor-dimension (e.g. $t=1_{S}$). Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is right adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$. In particular, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
@@
The desired claim follows if can find an isomorphism
@@
Note that Lemma˜3.2, as well as affineness being stable under base change, implies each vertical morphism in Equation˜3.1 are affine and locally of finite tor-dimension. Moreover, $(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}}\in D^{+}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ as upper shriek functors for proper morphisms preserve complexes with bounded below and coherent cohomology (see e.g. [Sta25, Tag 0AU1]). So, we have a string of isomorphisms
@@
By Lemma˜3.3, we see that $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor $\operatorname{Perf}(Y_{1}\times_{S}T)\to\operatorname{Perf}(Y_{2}\times_{S}T)$. Then [DLM25, Lemma 3.2] implies
@@
There is another string of isomorphisms:
@@
The last isomorphism comes from the work above, whereas the second isomorphism is [Riz17, Lemma 2.13], which completes the proof. ∎
@@
Let $f_{i}\colon Y_{i}\to S$ be proper flat morphisms to a Noetherian scheme where $i\in\{1,2\}$. Denote by $g_{i}\colon Y_{1}\times_{S}Y_{2}\to Y_{i}$ the natural morphisms. Suppose $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{1}$. Then $\Phi_{K}$ admits a left adjoint on $D_{\operatorname{qc}}$. In particular, the left adjoint is of the form $\Phi_{K^{\prime}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,g^{!}_{1}\mathcal{O}_{Y_{1}})$.
@@
$\operatorname{Hom}$ $(E,\Phi_{K}(G))\cong\operatorname{Hom}(E,\mathbf{R}(g_{2})_{\ast}(K\otimes^{\mathbf{L}}\mathbf{L}g_{1}^{\ast}G))$ (Definition)
$\cong\operatorname{Hom}(\mathbf{L}g_{2}^{\ast}E,K\otimes^{\mathbf{L}}\mathbf{L}g_{1}^{\ast}G)$ (Adjunction)
@@
Consider ˜3.1 where $t$ is an affine morphism locally of finite tor-dimension. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{1}$. Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is left adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})$. In particular, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.

### v2

@@ 1. Introduction
1.1. What is known
The context of our work starts with Mukai [Muk81]. Loc. cit. introduced integral transforms to study triangulated equivalences between derived categories associated with abelian varieties. Over time, integral transforms have played a central role in the interplay between derived categories and algebraic geometry (e.g. [Bri02, Kaw02]), with further applications in string theory [AHR05, ASS05].
Much of this progress has focused on smooth varieties, where geometric results rely on categorical methods. However, extending these methods to singular varieties remains challenging, largely because bounded coherent complexes need not be perfect. Thus, an important task is to develop techniques that improve our understanding of integral transforms in such singular settings.
@@ 1.1. What is known
1.2. What we do
In this work, we deepen this perspective by systematically exploiting special fibers. Specifically, we study the behavior of integral transforms under base change in settings that incorporate local algebra. Moreover, our approach naturally extends several results in the literature to arbitrary fields, and at the same time yields new insight for singularities of fibrations.
Preservation of adjointness
@@ 1.2. What we do
Consequently, questions about equivalences may be reduced to the behavior of integral transforms on these smaller subcategories. Whether an integral transform restricts to these smaller categories is now well-understood [Riz17, Bal09, DLM25]. These results shed light on the existence of left or right adjoints to integral transforms and the conditions under which such adjoints also restrict to these subcategories.
@@
Fully faithfulness & equivalences
With this understanding of adjointness under base change in hand, we turn to questions of fully faithfulness and equivalence. Particularly, regarding the preservation of adjointness and the restriction to smaller categories. Next, we shift gears to studying fully faithfulness and equivalences under base change; specifically, the ascent and descent of these properties. Before doing so, we review the relevant literature a bit more.
As mentioned earlier, the work of [HLS07, HLS09] studied these questions for fibrations whose base is of finite type over an algebraically closed field. A key ingredient in loc. cit. was to look at the behavior on special fibers. More recently, such questions have also been studied for smooth proper morphisms over a Noetherian base scheme (see [Kur24, Proposition/Definition 2.3]). However, smoothness assumptions restrict the range of examples one can consider, e.g. fibrations admitting singular fibers.
A result of Orlov addresses similar questions in the case of smooth varieties over a field. Specifically, for smooth projective varieties over a field, an integral transform is fully faithful or an equivalence if, and only if, the induced integral transform satisfies the same property after any base field extension [Orl02, Lemma 2.12]. The strategy of loc. cit. leverages ‘canonical morphisms’ between the kernels of integral transforms that describe the (co)units of adjunction. However, it remains unclear whether such canonical morphisms exist in the singular setting. This uncertainty provides further motivation for our work: we bypass the absence of such tools in the singular setting by developing different techniques.
This brings attention to the following, which summarizes Section˜4.
@@
1.3. Why it matters
@@
To conclude our work, we discuss the ramifications of the results above. Our results extend the literature on schemes of finite type over algebraically closed fields to the more general setting of arbitrary fields. They reveal a new interplay between the intuition of [HLS07, HLS09] concerning special fibers and faithfully flat descent along affine morphisms. In particular, certain singularities are shown to be derived invariants for fibrations over Noetherian schemes, broadening the scope to include arithmetically flavored situations of potential interest for future research.
@@
For a pair of geometrically integral projective curves over a field $k$ of characteristic zero, Fourier–Mukai partnership implies they are $k$-forms (see Proposition˜5.1). Here, a $k$-form means there exists a finite extension $L/k$ such that their base changes along $L/k$ become isomorphic. The result holds for singular curves and enriches algebraically closed case of [Spe23]. For fibrations of relative dimension one over a regular Noetherian $\mathbb{Q}$-scheme, Fourier–Mukai partnership forces the genera of special fibers to coincide (see Corollary˜5.2).
@@
Smoothness for proper schemes over an arbitrary field is likewise invariant under Fourier–Mukai partnership (see Proposition˜5.3). This subtlety arises especially over imperfect fields. Since regularity is characterized by $D^{b}_{\operatorname{coh}}=\operatorname{Perf}$, one might expect this condition to persist after a change of field extensions, but there exist regular projective curves that become singular after such a base change (see e.g. [Kol11, Remark 16]).
@@
In the setting of projective fibrations, smoothness of the total space is also a derived invariant if the base is regular (see Corollary˜5.4). Finally, generalizing [HLS09], we show that the Cohen–Macaulay and Gorenstein properties of a morphism remain invariant under Fourier–Mukai partnership for regular Noetherian base schemes (see Proposition˜5.5).
In a related but distinct direction, we study a variation of Proposition˜5.5 in the case of derived equivalences. Particularly, we prove a version for the Gorenstein condition for arbitrary derived equivalences where the required functor need not be an integral transform. See Proposition˜5.6.
@@
At times if $X$ is affine, we abuse notation and write $D_{\operatorname{qc}}(R):=D_{\operatorname{qc}}(X)$ where $R:=H^{0}(X,\mathcal{O}_{X})$ are the global sections; similar conventions will occur for the other categories mentioned here and later.
@@ 2.1. Perfect & pseudocoherent complexes
We say $P\in D_{\operatorname{qc}}(X)$ is perfect if it is affine locally isomorphic to a bounded complex of locally free sheaves of finite rank. Denote by $\operatorname{Perf}(X)$ for the full subcategory of perfect complexes. Also, given an integer $N$, we say $E\in D_{\operatorname{qc}}(X)$ is $N$-pseudocoherent if locally there is a morphism $P\to E$ from some $P\in\operatorname{Perf}(X)$ such that the induced morphism on cohomology sheaves $\mathcal{H}^{n}(P)\to\mathcal{H}^{n}(E)$ is an isomorphism for $n>N$ and is surjective for $n=N$. We say $E\in D_{\operatorname{qc}}(X)$ is pseudocoherent if it is $N$-pseudocoherent for every $N\in\mathbb{Z}$. In general, since $X$ is a scheme, $E\in D_{\operatorname{qc}}(X)$ is pseudocoherent if and only if locally on $X$ there is a quasi-isomorphism $P\to E$, with $P$ a bounded above complex of finite-rank free modules [GW23, Corollary 22.47]. As $X$ is Noetherian, $E\in D_{\operatorname{qc}}(X)$ is pseudocoherent if, and only if, $\mathcal{H}^{n}(E)$ is coherent for all $n$ and vanishes for $n\gg 0$ (see e.g. [Sta25, Tag 08E8]). The subcategory of pseudocoherent complexes of $D_{\operatorname{qc}}(X)$ coincides with $D^{-}_{\operatorname{coh}}(X)$ (i.e. the full subcategory of complexes with bounded above and coherent cohomology). Consider a morphism $f\colon Y\to X$ of Noetherian schemes. Then one has $\mathbf{L}f^{\ast}D^{-}_{\operatorname{coh}}(X)\subseteq D^{-}_{\operatorname{coh}}(Y)$ (see e.g. [Sta25, Tag 08H4]). We will use this characterization liberally.
@@ 2.3. Generation for triangulated categories
Let $\mathcal{T}$ be a triangulated category with shift functor $1]\colon\mathcal{T}\to\mathcal{T}$. Fix a subcategory $\mathcal{S}\subseteq\mathcal{T}$. Denote by $\operatorname{add}(\mathcal{S})$ the strictly full (that is, closed under isomorphisms) subcategory of $\mathcal{T}$ consisting of all direct summands of objects of the form $\oplus_{n\in\mathbb{Z}}S_{n}^{\oplus r_{n}}[n]$ where $S_{n}\in\mathcal{S}$ for all $n$ and finitely many $r_{n}$ are nonzero. Inductively, let
@@
Now, we impose $\mathcal{T}$ be a compactly generated triangulated category. To clarify, we require that $\mathcal{T}$ admits all small coproducts. Denote the collection of compact objects of $\mathcal{T}$ by $\mathcal{T}^{c}$. We say $G\in\mathcal{T}^{c}$ is a compact generator for $\mathcal{T}$ if for any $E\in\mathcal{T}$, one has $\operatorname{Hom}(G,E[n])=0$ for all $n\in\mathbb{Z}$ if, and only if, $E=0$. A well-known fact is that classical generators for $\mathcal{T}^{c}$ coincide with being compact generators for $\mathcal{T}$ (see e.g. [Sta25, Tag 09SR]).
@@
Let $X$ be a quasi-compact quasi-separated scheme. The collection of compact objects in $D_{\operatorname{qc}}(X)$ coincides with $\operatorname{Perf}(X)$. Moreover, $D_{\operatorname{qc}}(X)$ is compactly generated by a single object from $\operatorname{Perf}(X)$. See [BvdB03, Theorem 3.1.1] for details.
@@ 3. Preservation of adjointness
Consider ˜3.1 with $t$ a morphism locally of finite tor-dimension. Then $t_{1},t_{2},t^{\prime}$ are locally of finite tor-dimension. In particular, each face of Diagram˜3.1 is a tor-independent square (in the sense of [Lip09, Definition 3.10.2]).
@@
This first claim is affine local on $S$. Indeed, if needed, we can find an affine open cover of $S$ such that the base changes of $t$ along the associated open immersion are of finite tor-dimension. Then, for each such open immersion of a component in the affine cover, we have a similar diagram to Diagram˜3.1 (i.e. a ‘base change cube’). If we can show the desired properties on the morphisms arising in each diagram associated to the components of the affine open cover of $S$, then the desired claim follows. So, we may impose that $t$ be of finite tor-dimension.
By base change, we know that $g_{1},g_{2},f^{\prime}_{2},f^{\prime}_{1},g_{2}^{\prime},g_{1}^{\prime}$ are proper flat morphisms. Consider the fibered square in Diagram˜3.1 induced from the base change of $f_{1}$ along $t$. It follows by [GW23, Lemma 23.27] that $t_{1}$ is of finite tor-dimension. A similar argument shows $t_{2}$ is of finite tor-dimension as well. In fact, if we work with the fibered square in Diagram˜3.1 obtained by base change of $f_{2}^{\prime}$ along $t_{1}$, we see from similar reasoning that $t^{\prime}$ is locally of finite tor-dimension. This completes the proof for our first claim.
@@
Consider ˜3.1 with $t$ an affine morphism. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{1}$. Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor on $D^{b}_{\operatorname{coh}}$. Similarly, if $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{2}$, then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor on $\operatorname{Perf}$.
@@
We only prove the case of $K$ being relatively perfect over $Y_{2}$ implying $\Phi_{\mathbf{L}(t^{\prime})K}$ restricts to an exact functor on $\operatorname{Perf}$ because a similar argument applies for $D^{b}_{\operatorname{coh}}$. As $\Phi_{K}(\operatorname{Perf}(Y_{1}))\subseteq\operatorname{Perf}(Y_{2})$, [DLM25, Proposition 3.2] implies
@@
Consider ˜3.1. Let $K\in D_{\operatorname{qc}}(Y_{1}\times_{S}Y_{2})$. Then there is a natural isomorphism $\beta^{K}\colon\mathbf{L}t_{2}^{\ast}\circ\Phi_{K}\to\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}\circ\mathbf{L}t_{1}^{\ast}$ on $D_{\operatorname{qc}}$.
@@
This follows from the string of functorial isomorphisms for each $E\in D_{\operatorname{qc}}(Y_{1}\times_{S}T)$:
@@
However, we have an isomorphism in Diagram˜3.1,
@@
$=t_{2}(\operatorname{supp}(\mathcal{H}^{j}(\mathbf{L}t_{2}^{\ast}\Phi_{K}(E))))$ $\textrm{(Flatness of $t_{2}$)}.$
@@
Consider ˜3.1 where $t$ is an affine morphism locally of finite tor-dimension (e.g. $t=1_{S}$). Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{2}$. Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is right adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}})$. In this case, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.
@@
The desired claim follows if we can find an isomorphism
@@
Note that Lemma˜3.2, as well as affineness being stable under base change, implies each vertical morphism in Diagram˜3.1 are affine and locally of finite tor-dimension. Moreover, $(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{2}}\in D^{+}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ as upper shriek functors for proper morphisms preserve complexes with bounded below and coherent cohomology (see e.g. [Sta25, Tag 0AU1]). So, we have a string of isomorphisms
@@
in $D^{+}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2}\times_{S}T)$. By Lemma˜3.3, we see that $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ restricts to an exact functor $\operatorname{Perf}(Y_{1}\times_{S}T)\to\operatorname{Perf}(Y_{2}\times_{S}T)$. Then [DLM25, Lemma 3.2] implies
@@
There is another string of isomorphisms for all $A\in D_{\operatorname{qc}}$:
@@
The last isomorphism comes from the work above, whereas the second isomorphism is [Riz17, Lemma 2.13] together with the fact that $\mathbf{L}(t^{\prime})^{\ast}K$ is $g_{1}^{\prime}$-perfect (see see Corollary˜3.6 and the discussion in Section˜2.2), which completes the proof. ∎
@@
Let $f_{i}\colon Y_{i}\to S$ be proper flat morphisms to a Noetherian scheme where $i\in\{1,2\}$. Denote by $f^{\prime}_{i}\colon Y_{1}\times_{S}Y_{2}\to Y_{i}$ the natural morphisms. Suppose $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ is relatively perfect over $Y_{1}$. Then $\Phi_{K}$ admits a left adjoint on $D_{\operatorname{qc}}$. In particular, the left adjoint is of the form $\Phi_{K^{\prime}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{1})^{!}\mathcal{O}_{Y_{1}})$.
@@
$\operatorname{Hom}$ $(E,\Phi_{K}(G))\cong\operatorname{Hom}(E,\mathbf{R}(f_{2}^{\prime})_{\ast}(K\otimes^{\mathbf{L}}\mathbf{L}(f^{\prime}_{1})^{\ast}G))$ (Definition)
$\cong\operatorname{Hom}(\mathbf{L}(f_{2}^{\prime})^{\ast}E,K\otimes^{\mathbf{L}}\mathbf{L}(f_{1}^{\prime})^{\ast}G)$ (Adjunction)
@@
Consider ˜3.1 where $t$ is an affine morphism locally of finite tor-dimension. Let $K\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$ be relatively perfect over $Y_{1}$. Then $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K^{\prime}}$ is left adjoint to $\Phi_{\mathbf{L}(t^{\prime})^{\ast}K}$ on $D_{\operatorname{qc}}$ where $K^{\prime}:=\mathbf{R}\operatorname{\mathcal{H}\!\mathit{om}}(K,(f^{\prime}_{2})^{!}\mathcal{O}_{Y_{1}})$. In such a case, $K^{\prime}\in D^{b}_{\operatorname{coh}}(Y_{1}\times_{S}Y_{2})$.